A ball is thrown vertically upward from ground level hits the ground after 4s.calculate the maximum height it reached during it's journey

1
Expert's answer
2021-11-09T10:59:00-0500


The maximum height is



"H=\frac{g(T/2)^2}{2}=\frac{gT^2}{8}=\frac{9.8·4^2}{8}=19.6\text{ m}."